OpenStudy (anonymous):

Use the function to answer the question. f(x) = -2|x - 5| + 3 What is the vertex of the graph of f(x)? (Points : 4) (-5, 3) (-2, 3) (5, -3) (5, 3)

OpenStudy (anonymous):

there is a rule with l x - 5l i know i wont have a negative

OpenStudy (cwrw238):

try drawing the graph

OpenStudy (cwrw238):



OpenStudy (cwrw238):

the 'V' is the graph of y = |x - 5| now can u continue?

OpenStudy (anonymous):

im still lost here

OpenStudy (cwrw238):

the |x-5| is obtained by defecting y = x-5 in the x axis (all values of y are positive

OpenStudy (anonymous):

is it supposed to be 5,0?

OpenStudy (cwrw238):

no you now have to transform y = |x - 5| to y = -2|x - 5| + 3 the negative value will reflect the whole graph in x axis and the + 3 will raise the whole graph by 3 units in y direction

OpenStudy (anonymous):

so the v will be faced down and moved up 3 on the y axis?

OpenStudy (anonymous):

so it would be 5,3

OpenStudy (cwrw238):

i'm a bit rusty with these - hold on a sec

OpenStudy (anonymous):

ok :)

OpenStudy (jdoe0001):

look at the absolute value expression, what value of "x" makes |x - 5| = 0?

OpenStudy (cwrw238):

cant remember jdoe will solve it

OpenStudy (anonymous):

5-5 = 0

OpenStudy (jdoe0001):

then \(\bf -2|x - 5| + 3\\ \quad \\ -2|(\color{red}{5}) - 5| \color{red}{+ 3}\\ \quad \\ (\color{red}{5\quad ,\quad 3})\Leftarrow vertex\)

OpenStudy (anonymous):

ok i see thank you ^^

OpenStudy (jdoe0001):

yw

OpenStudy (cwrw238):

of course - i had a mental aberation !!!